OpenStudy (anonymous):

A circle has an area twice that of a square. Find the ratio of the circumference of the circle to the perimeter of the square.

OpenStudy (amistre64):

squaring the circle are we ..

OpenStudy (amistre64):

so a ratio of perimeters ....

OpenStudy (anonymous):

I don't even know where to start. I mean, I could obviously guess and check, but I wanted to know how to do it algebraically.

OpenStudy (amistre64):

Sa = s^2 Ca = pi r^2 = 2Sa = 2s^2 so far

OpenStudy (anonymous):

25a?

OpenStudy (anonymous):

Oh, 2Sa

OpenStudy (anonymous):

sorry

OpenStudy (amistre64):

4s = perimeter square 2pi r = perimeter of a circle pi r^2 = 2s^2 is a good relationship to figure out the rest

OpenStudy (amistre64):

prolly get things in terms of s

OpenStudy (anonymous):

sqrt(pir^2/2) = s

OpenStudy (anonymous):

r = sqrt(2s^2/pi)

OpenStudy (anonymous):

where from there?

OpenStudy (amistre64):

so depending on what variable youwant this to speak in: 4s : 2pi r and then simplify is my best guess

OpenStudy (anonymous):

apply them to the perimeter equations then square it?

OpenStudy (amistre64):

a ratio is just how one thing looks next to another; proportions and fractions and the like

OpenStudy (anonymous):

yes 2pisqrt(2s^2/pi) : 4sqrt(pir^2/2)

OpenStudy (anonymous):

ima square both sides

OpenStudy (amistre64):

get them speaking the same variable; not 2 different variables

OpenStudy (anonymous):

how?

OpenStudy (amistre64):

by solving for s or r from the relationship itself; pick on of them and go with it :)

OpenStudy (anonymous):

but all we have so far is the area and perimeter formulae

OpenStudy (amistre64):

and how r and s are related pi r^2 = 2s^2

OpenStudy (amistre64):

the area of the circles is two times that of the area of the square

OpenStudy (anonymous):

you're gonna have to do all of this I'm not following

OpenStudy (amistre64):

ok, heres my thoughts area square = s^2 area circle = pi r^2 right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

area circle = 2 times area square pi r^2 = 2s^2 right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

good, solve for say ... r in terms of s

OpenStudy (anonymous):

r = sqrt(2s^2/pi)

OpenStudy (amistre64):

pi r^2 = 2s^2 r^2 = 2s^2 /pi r = s sqrt(2/pi)

OpenStudy (amistre64):

now; perimeter of a square is: 4s perimeter of a circle is: 2pi r right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

so the ratio of circle to square is then 4s : 2pi r or simply s = pi r now sub in for r

OpenStudy (amistre64):

not s = ... but typos persist

OpenStudy (anonymous):

?

OpenStudy (anonymous):

s = pi r?

OpenStudy (anonymous):

wouldnt it be 2s = pi r?

OpenStudy (amistre64):

so the ratio of circle to square is then 4s : 2pi r or simply s : pi r now sub in for r ; we know r = s sqrt(2/pi) s : s pi sqrt(2/pi) divide off the s 1 : pi sqrt(2/pi)

OpenStudy (amistre64):

like i said, typos persist :)

OpenStudy (anonymous):

what do you mean by that? sorry I was away for a couple of minutes, the site was doing weird stuff

OpenStudy (amistre64):

yeah, its freezing up on my end too

OpenStudy (anonymous):

"typos persist" whhat do you mean by that?

OpenStudy (amistre64):

i mean that you have to read thru the errors that pop up from trying to type on a keyboard ...

OpenStudy (anonymous):

so will you please fix that whole post, (your fourth to last)

OpenStudy (amistre64):

so the ratio of circle to square is then 4s : 2pi r or simply s : pi r now sub in for r ; we know r = s sqrt(2/pi) s : s pi sqrt(2/pi) divide off the s 1 : pi sqrt(2/pi) we can push the pi inside by ^2ing it 1 : sqrt(2pi^2/pi) 1 : sqrt(2 pi) and thats about it

OpenStudy (anonymous):

s : pi r you did it again

OpenStudy (anonymous):

alright, I still dont understand what your typo problem was... however you did get the answer for your typo, and Now I can just do sqrt(2pi) : 2

OpenStudy (amistre64):

lol .... yeah :)

OpenStudy (amistre64):

4/2 = 2 not 1 ... i really should try to sleep ;)

OpenStudy (anonymous):

That's probably a good idea